# Detecting Change Points of Covariance Matrices in High Dimensions

### v1

@@ A.3 Auxiliary results
Next, we prove Lemma 4 providing an approximation for the quantity $\sigma_{1}^{2}$ defined below (4.23).
@@
×$\quad\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)$ (A.24)
@@
for $1\leq i\leq p,~{}1\leq j<k\leq p.$ Then, we may write (replacing for a moment $i$ by $i-1$)
@@
Calculation of $S_{i,1}$
By an application of the Sherman-Morrison formula, we obtain
@@
Applying the trace on both sides and dividing by $k-j+1$, yields
@@
Moreover, note that $||\mathbf{S}_{i,j:k}^{-1}||\leq 1/((1-\sqrt{t_{0}})^{2}-\varepsilon)<\infty$ for some $\varepsilon>0$ and all large $n.$ As a further preparation, we note that $(\lfloor nt_{1}\rfloor/n)*(1/i)\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$ can be approximated by the first negative moment of the Marčenko–Pastur distribution $F^{i/\lfloor nt_{2}\rfloor}$, that is,
@@
Calculation of $S_{i,2}$
Similarly to the previous step, we may show that
@@
Calculation of $\mathbf{S}_{i,3,1}$
Applying similar techniques as in the previous steps, we get
@@
Calculation of $\mathbf{S}_{i,3,2}$
Again applying similar techniques as in the previous steps, especially (A.26), (A.27) and (A.31), we get
@@
×$\quad~{}\quad~{}\quad~{}\quad~{}\quad~{}\quad\times\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\Big{\}}$
@@
×$\quad\times\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
@@
×$\quad\times\Big{\{}1+\frac{i^{2}}{\left(n-\lfloor nt_{1}\rfloor\right)\left(\lfloor nt_{2}\rfloor\right)}-\frac{i}{n-\lfloor nt_{1}\rfloor}-\frac{i}{\lfloor nt_{2}\rfloor}\Big{\}}+o_{\mathbb{P}}(1)$
@@
Conclusion
Using (A.24) and (A.25), we obtain
@@ Conclusion
[…]
@@
For the proof of Theorem 4, we approximated the quadratic term appearing in the decomposition of $\log\Lambda_{n,t}$ by a deterministic term. To analyze its asymptotic behaviour, we first state the consistency of the kurtosis estimator, which will be proved in Section A.4.
Lemma 9.
Suppose that assumptions (A1)-(A3) are satisfied, and that $H_{0}$ holds true. Then,
$\frac{\hat{\kappa}_{n}}{\kappa_{n}}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}1.$
@@
where we used Lemma 9. This implies assertion (A.42). Thus, it remains to show that
@@
which follows from (4.27) and (A.64) given later. ∎
A.4 Proof of Lemma 9
@@
$\kappa=3+\frac{\nu_{n}-2\tau_{n}}{\omega_{n}}.$
Following the routine in Section S1$\cdot$1 of Lopes, Blandino and Aue (2019), the assertion of Lemma 9 is implied by the following results.
@@
Suppose that assumptions (A1) and (A2) are satisfied, and that $H_{0}$ holds true. Then, it holds that
@@
Then, it follows from Lemma S.2 in Lopes, Blandino and Aue (2019) that
@@
We continue with studying the second term $\hat{\omega}_{n,2}.$ Without loss of generality, we may assume that $\mathbb{E}[y_{j1}]=0$ for all $1\leq j\leq p$, and we use the notation $(U_{kl})_{1\leq k,l\leq p}=\mathbf{\Sigma}^{1/2}.$ As a preparation, we note that $\mathbb{E}[y_{j1}^{2}]=\Sigma_{jj}\leq\|\mathbf{\Sigma}\|\lesssim 1$. Moreover, note that $\max_{1\leq k,l\leq p}|U_{kl}|\lesssim 1$, where $U_{kl}$ denote the entries of $\mathbf{\Sigma}^{1/2}$. Then, one can also verify by a direct calculation $\mathbb{E}[y_{j1}^{4}]\lesssim 1$. These considerations imply
@@
For a proof of part (c) we note Lemma S.3 in Lopes, Blandino and Aue (2019) implies
@@
Step 1
Let $\overline{y}_{1\cdot},\ldots,\overline{y}_{p\cdot}$ denote the components of the $p$-dimensional vector $\overline{\mathbf{y}}.$ Then, a direct computation gives
@@
In summary, we obtain using (A.50), (A.52), (A.58) and assumption (A3)
$\mathbb{E}(\breve{\nu}_{n,1}^{1/2}-\hat{\nu}_{n,1}^{1/2})^{2}\lesssim 1+\frac{1}{n}||\mathbf{\Sigma}||_{F}^{2}\lesssim 1.$ (A.59)
Step 2
Note that $\breve{\nu}_{n}$ is unbiased for $\nu_{n}.$ Therefore, we get
@@
Conclusion
Using (A.59), (A.60), (A.61) and $\nu_{n}\gtrsim n$ (see p.3 in the supplementary material of Lopes, Blandino and Aue, 2019), we obtain
@@ A.5 Proofs of Lemma 5 - Lemma 6
Note that the proof of Lemma 2 is very similar to the proof of Lemma 3 in Dörnemann (2023) and we skip it for the sake of brevity.
@@
where $A_{it}$ is defined in (A.41). Then, the decomposition (4.26) is obviously true. Note that the definition of $\breve{\sigma}_{n,t}^{2}$ in (4.17) implies
@@
almost surely. Thus, we conclude that $(Q_{n,1,t})$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}])$, and it remains to show (4.27). Applying Lemma 2.2 in Li (2003), we obtain

### v2

@@ A.3 Auxiliary results
In this section, we provide the proofs of the auxiliary results given in section A.1, among others. Note that the proof of Lemma 2 is very similar to the proof of Lemma 3 in Dörnemann (2023) and we skip it for the sake of brevity. To begin with, we prove Lemma 4 providing an approximation for the quantity $\sigma_{1}^{2}$ defined below (4.27).
@@
$\quad\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;(\lfloor nt_{1}\rfloor+1):n)\mathbf{P}^{(\lfloor nt_{1}\rfloor+1):\lfloor nt_{2}\rfloor}(i-1;1:\lfloor nt_{2}\rfloor)\right)$ (A.7)
@@
for $1\leq i\leq p,~1\leq j<k\leq p.$ Then, we may write (replacing for a moment $i$ by i1i-1)
@@
Calculation of $S_{i,1}$ By an application of the Sherman-Morrison formula, we obtain
@@
Applying the trace on both sides and dividing by kj+1k-j+1, yields
@@
Moreover, note that $\mathbf{S}_{i,j:k}^{-1}||\leq 1/((1-\sqrt{t_{0}})^{2}-\varepsilon)<\infty$ for some $\varepsilon>0$ and all large $n.$ As a further preparation, we note that $(\lfloor nt_{1}\rfloor/n)*(1/i)\operatorname{tr}\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{-1}$ can be approximated by the first negative moment of the Marčenko–Pastur distribution $F^{i/\lfloor nt_{2}\rfloor}$, that is,
@@
Calculation of $S_{i,2}$ Similarly to the previous step, we may show that
@@
Calculation of $\mathbf{S}_{i,3,1}$ Applying similar techniques as in the previous steps, we get
@@
Calculation of $\mathbf{S}_{i,3,2}$ Again applying similar techniques as in the previous steps, especially (A.9), (A.10) and (A.14), we get
@@
$\quad~\quad~\quad~\quad~\quad~\quad\times\mathbf{x}_{i,l}^{\top}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}^{(-l)}\right)^{-1}\mathbf{x}_{i,k}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}^{(-l)}\right)^{-1}\mathbf{x}_{i,l}\Big\}$
@@
$\quad\times\sum_{\begin{subarray}{c}k=\lfloor nt_{1}\rfloor+1\end{subarray}}^{\lfloor nt_{2}\rfloor}\mathbf{x}_{i,k}^{\top}\left(\mathbf{S}_{i,(\lfloor nt_{1}\rfloor+1):n}\right)^{-1}\left(\mathbf{S}_{i,1:\lfloor nt_{2}\rfloor}\right)^{-1}\mathbf{x}_{i,k}$
@@
$\quad\times\Big\{1+\frac{i^{2}}{\left(n-\lfloor nt_{1}\rfloor\right)\left(\lfloor nt_{2}\rfloor\right)}-\frac{i}{n-\lfloor nt_{1}\rfloor}-\frac{i}{\lfloor nt_{2}\rfloor}\Big\}+o_{\mathbb{P}}(1)$
@@
Conclusion Using (A.7) and (A.8), we obtain
@@
$=2(t_{2}-t_{1})^{2}\int_{0}^{1}\frac{yx}{y^{-1}(1-t_{1})t_{2}-(t_{2}-t_{1})x}dx+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\int_{0}^{1}\frac{y^{2}x}{(1-t_{1})t_{2}-(t_{2}-t_{1})yx}dx+o_{\mathbb{P}}(1)$
$=2(t_{2}-t_{1})^{2}\int_{0}^{y}\frac{x}{(1-t_{1})t_{2}-(t_{2}-t_{1})x}dx+o_{\mathbb{P}}(1)$
@@
where we used Proposition 1. This implies assertion (A.25). Thus, it remains to show that
@@
which follows from (A.2) and (A.47) given later. ∎
A.4 Proof of Proposition 1
@@
$\kappa_{n}=3+\frac{\nu_{n}-2\tau_{n}}{\omega_{n}}.$
[…]
@@
Suppose that assumptions (A-1) and (A-2) are satisfied, and that $H_{0}$ holds true. Then, it holds that
@@
Then, it follows from Lemma S.2 in Lopes et al. (2019) that
@@
We continue with studying the second term $\hat{\omega}_{n,2}.$ Without loss of generality, we may assume that $\mathbb{E}[y_{j1}]=0$ for all $1\leq j\leq p$, and we use the notation $(U_{kl})_{1\leq k,l\leq p}=\boldsymbol{\Sigma}^{1/2}.$ As a preparation, we note that $\mathbb{E}[y_{j1}^{2}]=\Sigma_{jj}\leq\|\boldsymbol{\Sigma}\|\lesssim 1$. Moreover, note that $\max_{1\leq k,l\leq p}|U_{kl}|\lesssim 1$, where $U_{kl}$ denote the entries of $\boldsymbol{\Sigma}^{1/2}$. Then, one can also verify by a direct calculation $\mathbb{E}[y_{j1}^{4}]\lesssim 1$. These considerations imply
@@
For a proof of part (c) we note Lemma S.3 in Lopes et al. (2019) implies
@@
Step 1 Let $\overline{y}_{1\cdot},\ldots,\overline{y}_{p\cdot}$ denote the components of the $p$-dimensional vector $\overline{\mathbf{y}}.$ Then, a direct computation gives
@@
In summary, we obtain using (A.33), (A.35), (A.41) and assumption (A-3)
[…]
Step 2 Note that $\breve{\nu}_{n}$ is unbiased for $\nu_{n}.$ Therefore, we get
@@
Conclusion Using (A.42), (A.43), (A.44) and $\nu_{n}\gtrsim n$ (see p.3 in the supplementary material of Lopes et al., 2019), we obtain
@@ A.5 Proofs of Lemma 5 - Lemma 6
where $A_{it}$ is defined in (A.24). Then, the decomposition (A.1) is obviously true. Note that the definition of $\breve{\sigma}_{n,t}^{2}$ in (4.20) implies
@@
almost surely. Thus, we conclude that $(Q_{n,1,t})$ is asymptotically tight in the space $\ell^{\infty}([t_{0},1-t_{0}])$, and it remains to show (A.2). Applying Lemma 2.2 in Li (2003), we obtain